OpenStudy (anonymous):

find all theta in the interval 0 2pi that satisfy sin2theta=0

OpenStudy (anonymous):

\[\sin n\pi=0\] for \[n=0,1,2,...\]so \[\theta=\frac{n\pi}{2}\]